# Aligning Language Models with Real-time Knowledge Editing

### v2

@@ Abstract
Knowledge editing aims to modify outdated knowledge in large language models (LLMs) efficiently while retaining their original capabilities. Mainstream benchmarks for knowledge editing are predominantly static and fail to keep in pace with the evolving real-world knowledge. In this work, we introduce CRAFT, an ever-evolving real-world benchmark for knowledge editing. It features well-designed paired edits for composite reasoning, and evaluates models on alias portability as well as temporal and common-sense locality, making it a challenging knowledge editing benchmark on which previous knowledge editing methods hardly achieve balanced performance. Towards flexible real-time editing, we propose KEDAS, a novel paradigm of knowledge editing alignment featuring diverse edit augmentation and self-adaptive post-alignment inference, which exhibits significant performance gain on CRAFT compared to previous methods. All of our code and data are available at https://anonymous.4open.science/r/CRAFT-KEDAS.
@@
Chenming Tang   Yutong Yang   Kexue Wang   Yunfang Wuthanks: Corresponding author. National Key Laboratory for Multimedia Information Processing, Peking University MOE Key Laboratory of Computational Linguistics, Peking University School of Computer Science, Peking University {tangchenming, 2300013219}@stu.pku.edu.cn {yytpku, wuyf}@pku.edu.cn
$\dagger$$\dagger$}\$footnotetext: These authors contributed equally.
@@ 1 Introduction
Large language models (LLMs) Llama Team (2024); Gemini Team (2025) have been the core of modern natural language processing (NLP). Once pre-trained, knowledge is injected into LLMs and becomes their static internal capability Petroni et al. (2019). As time goes by, some knowledge in LLMs inevitably becomes incorrect or out of date, and it is vital to update the outdated knowledge. However, retraining a model from scratch is highly costly, especially for downstream users. To this end, the techniques of knowledge editing have been developed Zhang et al. (2024), aimed at editing specific knowledge in LLMs efficiently without heavy re-training.
There have been a large number of benchmarks to evaluate the effectiveness of knowledge editing methods. Nevertheless, most traditional knowledge editing benchmarks like ZsRE Levy et al. (2017), MQuAKE Zhong et al. (2023) and RippleEdits Cohen et al. (2024a) are static and not real-time. Once the datasets are constructed and released, they remain fixed and can not be updated anymore, which is far from real-world knowledge editing applications. Recently, some work is devoted to building real-time datasets. For example, WikiBigEdit Thede et al. (2025) collects real-world changes in Wikipedia, but it requires processing massive Wiki data (usually hundreds of gigabytes) from the Internet, which is costly and inconvenient in practice. Meanwhile, the automatically collected dataset lacks necessary filtering and thus suffers from severe sparsity (for example, only 3.6% of the collected data have an entry of portability evaluation). EvoWiki Tang et al. (2025) is another real-time dataset also based on Wikipedia, but it focuses on retrieval-augmented generation (RAG) and continual learning. Overall, there still lacks a high-quality, easily-updated benchmark for real-world and real-time knowledge editing.
Benchmark Real- time Real- world Reasoning Portability Composite Reasoning Portability
@@
Table 1: Comparison of knowledge editing benchmarks. Composite reasoning portability evaluates a model’s ability to integrate multiple distinct edits into a single reasoning query, whereas reasoning portability tests reasoning chains derived from a single edit.
In this work, we introduce CRAFT (Chinese Real-time statistics And Finance knowledge ediTing benchmark), an ever-evolving real-world benchmark for knowledge editing in Chinese. It leverages publicly available official data that are continuously updated, ensuring both temporal freshness and real-world applications. Furthermore, it is organized in well-designed paired edits, each serving as a composite reasoning test to evaluate models’ ability to integrate multiple related factual updates. Moreover, CRAFT supports evaluations on alias portability and temporal and common-sense locality, providing a comprehensive assessment of model adaptability and robustness under dynamic knowledge changes. The differences between CRAFT and previous datasets are summarized in Table 1.
We assess the exposure rate of CRAFT and existing benchmarks with five different LLMs, either open-source or closed-source. The results demonstrate that a large portion of knowledge in existing benchmarks has been known to LLMs while the exposure rate of CRAFT is nearly zero on all the five models. This reveals the disadvantages of static datasets and validates the real-time property of CRAFT, which makes it challenging to LLMs.
We evaluate a suite of representative knowledge editing methods on CRAFT and demonstrate that current approaches exhibit inherent limitations. For example, parameter-based approaches like ROME Meng et al. (2022) and WISE Wang et al. (2024a) suffer from gradual model degradation and struggle to achieve successful editing in the setting of sequential editing. Simple retrieval-based approaches like IKE Zheng et al. (2023) and EREN Chen et al. (2024b) fail to achieve consistent balanced performance due to lack of alignment. LTE Jiang et al. (2024), which aligns LLMs with knowledge editing via post-training, shows weak locality due to overfitting on irrelevant queries.
@@
  • We introduce CRAFT, a high-quality, easily-updated and fully open-source benchmark for real-time knowledge editing.
  • We evaluate existing knowledge editing methods on CRAFT and reveal the limitations of these methods in real-time knowledge editing.
  • We propose KEDAS, a novel paradigm of aligning LLMs with real-time knowledge editing featuring diverse edit augmentation and self-adaptive post-alignment inference, enabling dynamic routing between pre- and post-alignment LLMs, which significantly outperforms existing methods on CRAFT.
@@ 2.1 Task Formulation
An LLM system $f$ can be regarded as a function $f:\mathcal{Q}\mapsto\mathcal{A}$ mapping an query $q\in\mathcal{Q}$ to its output answer $a=f(q)\in\mathcal{A}$.
Knowledge editing aims to change the behavior of the LLM after modifying some knowledge. Edits of knowledge are usually in the form of query-answer (QA) pairs:
@@
To assess the efficacy of editing, the post-edit LLM $f^{*}$ is evaluated via the following criteria Zhang et al. (2024):
@@ 2.2 Settings of Knowledge Editing
This traditional setting is widely employed by earlier work. It is far from real-world applications, since only one edit can be applied each time.
@@
In sequential editing, edits are applied step by step (let $f^{0}$ denote $f$ for consistency):
@@
Knowledge Editing Alignment
LTE Jiang et al. (2024) relates knowledge editing to LLM alignment, aligning LLMs with ever-changing real-time knowledge edits. During alignment, LLMs’ capabilities of knowledge updating are enlightened by an editing prompt:
@@
In the inference phase, LLMs are required to conduct on-the-fly and streaming knowledge editing by retrieving relevant edits to the query from the stored memory.
@@
MQuAKE Zhong et al. (2023) 35.46% 76.21% 46.11% 42.58% 76.09% 5,604
EvoWiki_evolved Tang et al. (2025) 8.07% 16.22% 6.43% 10.24% 5.08% 1,338
EvoWiki_stable Tang et al. (2025) 28.45% 42.37% 25.90% 39.76% 12.32% 1,494
EvoWiki_uncharted Tang et al. (2025) 8.19% 22.62% 7.39% 16.02% 3.43% 1,136
@@ 3 The CRAFT Dataset
We introduce CRAFT, a real-time, real-world benchmark designed to evaluate knowledge editing under dynamically evolving factual contexts. Unlike static benchmarks that rely on outdated or widely exposed Wikipedia facts, CRAFT continuously updates with publicly available official data sources, ensuring fairness and real-world relevance.
3.1 Dataset Overview
CRAFT consists of two complementary subsets: CRAFT-Statistics and CRAFT-Finance, representing stable and highly correlated factual domains that evolve periodically. Example instances from the two subsets are shown in Appendix E and F respectively.
@@ 3.1 Dataset Overview
We collect monthly statistical reports from the official National Bureau of Statistics of China 222 https://data.stats.gov.cn/, covering July 2024 to June 2025 in the current version used in this work. The dataset spans four major domains—finance, fiscal affairs, telecommunications, and transportation—containing 221 indicators and 4,468 data points. Our data collection pipeline, leveraging the cn-stats API 333 https://github.com/songjian/cnstats, supports automatic temporal updates, enabling effortless retrieval of the most recent monthly data.
@@
CRAFT-Finance includes annual financial statements of 390 publicly listed Chinese companies (2023–2024), sourced from Eastmoney Financial Database 444 https://data.eastmoney.com/ via the AKShare API 555 https://github.com/akfamily/akshare. It features financial indicators such as debt ratio, shareholder equity, and operating cash flow, supporting yearly updates to reflect the latest financial disclosures.
@@ 3.2 Evaluation Metrics
To comprehensively assess model adaptability, besides the conventional edit success (ES) as described in Section 2.1, CRAFT provides four evaluation dimensions:
@@
$q_{p}$ $=\text{2024 Company A's debt ratio}$
$a_{p}$ $=0.5$
CRAFT-Finance includes ten types of such combinations, demonstrating the generality and practical relevance of composite reasoning portability. For CRAFT-Statistics, composite portability is constructed by pairing two edits of the same indicator from adjacent years.
@@ 3.3 Dataset Statistics
Table 2 compares the exposure ratio of CRAFT and existing benchmarks across five representative LLMs. Due to its temporal freshness and domain specificity, CRAFT exhibits extremely low exposure rates across all the five LLMs, mitigating data contamination and ensuring fair evaluation. Meanwhile, for other datasets, a great portion of the knowledge has been known to LLMs, making them less challenging and less fair in practice.
Dataset Edit $\mathbf{P}_{\text{reasoning}}$ $\mathbf{P}_{\text{alias}}$ $\mathbf{L}_{\text{temporal}}$ $\mathbf{L}_{\text{common}}$
@@
Table 3: Number of instances per evaluation type in CRAFT. Composite reasoning uses edit pairs; other metrics are per edit.
Table 3 shows instance counts in CRAFT sub-datasets. Composite reasoning portability is evaluated per edit pair, and other metrics per edit. For experiments, we split CRAFT into training and test sets. The test set contains 500 instances from the two subsets each (1,000 instances in total) and the training set contains the remaining instances.
3.4 Data Summary
CRAFT offers a scalable and continuously evolving benchmark reflecting real-world knowledge updates. Its dynamic nature, factual authenticity, and comprehensive reasoning evaluation make it a robust foundation for future research on reliable and interpretable knowledge editing. Anyone who wants to construct CRAFT within an arbitrary time period can easily obtain a customized benchmark with our provided open-source scripts.
@@ 4 KEDAS: A Proposal for Real-time Knowledge Editing Alignment
As will be shown in the experiments, previous methods show certain limitations either in edit success and portability or locality on CRAFT. To improve edit success and generalization while keeping a good locality in real-time knowledge editing, we propose KEDAS, an all-round approach consisting of several complementary techniques.
The overall framework of KEDAS is illustrated in Figure 1, including alignment, editing and inference. In alignment, the LLM is aligned with knowledge editing using LoRA. In editing, each edit is converted into diverse forms and stored to the memory module. In the inference phase, edits are retrieved from the memory based on the user query and then filtered by a trained classifier. If there exists relevant edits, the model with trained LoRA adapters will be called for inference. Otherwise, the query will go through the frozen base model. The order of editing and inference is flexible based on practical needs (“edit as you go"), as shown by the bidirectional arrow. For example, we can apply several edits and employ the LLM for inference immediately. If there come further new edits, we can return to the editing phase to conduct these edits and then resume inference.
@@ 4.2 Diverse Edit Augmentation
In previous work, each piece of edited knowledge is directly applied to the system as it is, limiting generalization. Since each knowledge can be expressed in various ways, leveraging just one fixed form is inflexible, especially for challenging portability evaluations like in CRAFT.
@@
Concretely, for an edit $e^{t}=(q^{t}_{e},a^{t}_{e})$ at step $t$ 666 For simplicity, we deem the two edits in each CRAFT instance as two separate time steps., its conventional QA form is:
@@
Specifically, declarative is to convert the QA pair into a declaration while aliased is to replace specialized terms in the declaration with their aliases for more flexible retrieval.
In experiments, we conduct these augmentations by prompting gpt-4o-mini 777 https://platform.openai.com, an API-based off-the-shelf LLM. The prompts are in Appendix C.
@@ 4.3 Filter-based Smart Retriever
During inference, the common practice regarding retrieval in previous knowledge editing work is to directly employ an embedder and perform single-stage retrieval to obtain the top-$k$ hits. This suffers from a poor locality due to the lack of a filter measuring the relevance of queries. In this work, we propose a filter-based smart retriever that pairs the high-recall retrieval with a high-precision filter by precisely identifying relevant queries. In this way, on challenging benchmarks like CRAFT, the post-edit system can keep consistent behaviors for neighboring locality queries while preserving its answers to general queries.
Firstly, we retrieve top-$n$ candidates from the memory using a normal embedding-based retriever. Thanks to our diverse edit augmentation, the memory contains various forms of edits, improving the recall rate. At the time step $t$, for an input query $q$, a set of top-$n$ candidates $\mathcal{C}^{\text{top-}n}$ is retrieved as:
@@
$\mathcal{E}^{*}=\operatorname{arg\,top}_{k}(\mathbf{Similarity}(e,q))\ \text{for }e\in\mathcal{C}^{\text{filtered}}.$ (13)
The filter is trained based on the training set of CRAFT. Details are specified in Appendix D.

### v3

@@ Abstract
Knowledge editing aims to modify outdated knowledge in language models efficiently while retaining their original capabilities. Mainstream datasets for knowledge editing are predominantly static and fail to keep in pace with the evolving real-world knowledge. In this work, we introduce CRAFT, an ever-evolving real-world dataset for knowledge editing. It evaluates models on temporal locality, common-sense locality, composite portability and alias portability, providing a comprehensive and challenging evaluation for knowledge editing, on which previous methods hardly achieve balanced performance. Towards flexible real-time knowledge editing, we propose KEDAS, a novel paradigm of knowledge editing alignment featuring diverse edit augmentation and self-adaptive post-alignment inference, exhibiting significant performance gain on both CRAFT and traditional datasets compared to previous methods. We hope this work may serve as a catalyst for shifting the focus of knowledge editing from static update to dynamic evolution. [^1] All of the code and data of this work are publicly available at https://github.com/JamyDon/CRAFT-KEDAS.
@@
Chenming Tang   Yutong Yang   Kexue Wang   Yunfang Wu National Key Laboratory for Multimedia Information Processing, Peking University School of Computer Science, Peking University {tangchenming, 2300013219}@stu.pku.edu.cn  {yytpku, wuyf}@pku.edu.cn Equal contribution  Corresponding author
@@ 1 Introduction
Once language models (LMs) are pre-trained, knowledge is injected into them and becomes their static internal capability Petroni et al. (2019). As time goes by, some knowledge in LMs inevitably becomes incorrect or out of date, and it is vital to update the outdated knowledge. However, re-training a model frequently is highly costly and usually infeasible. To this end, the techniques of knowledge editing have been developed Zhang et al. (2024) to edit specific knowledge in LMs efficiently without heavy re-training.
There have been a large number of datasets to evaluate the effectiveness of knowledge editing approaches. However, most traditional ones like ZsRE Levy et al. (2017), MQuAKE Zhong et al. (2023) and RippleEdits Cohen et al. (2024a) are static and not real-time. Once the datasets are constructed and released, they remain fixed and can not be updated anymore, which is far from real-world knowledge editing applications. Recently, there have been a few real-time datasets released. For example, WikiBigEdit Thede et al. (2025) collects real-world changes in Wikipedia, but it requires processing massive Wiki data (hundreds of gigabytes) from the Internet, which is costly and inconvenient in practice. Meanwhile, the automatically collected dataset lacks necessary filtering and thus suffers from severe sparsity (for instance, only 3.6% of the collected data have an entry of portability evaluation). EvoWiki Tang et al. (2025) is another real-time dataset based on Wikipedia, but it focuses on retrieval-augmented generation (RAG) and continual learning and does not support comprehensive evaluation for knowledge editing due to lack of crucial aspects like locality. Overall, there still lacks a high-quality, easily-updated dataset for real-world and real-time knowledge editing.
Dataset RT RW RP CRP
@@
Table 1: Comparison of knowledge editing datasets. RT refers to “real-time" and RW denotes “real-world". RP (reasoning portability) tests reasoning chains derived from a single edit, whereas CRP (composite reasoning portability) evaluates a model’s ability to integrate multiple distinct edits into a single reasoning query.
In this work, we introduce CRAFT (Chinese Real-time statistics And Finance knowledge editing dataseT), an ever-evolving real-world dataset for knowledge editing in Chinese. It leverages publicly available official data that are continuously updated, ensuring both temporal freshness and real-world applications. CRAFT is organized in well-designed paired edits, each serving as a composite reasoning test to evaluate models’ ability to integrate multiple related factual updates. Meanwhile, it supports evaluations on alias portability and temporal and common-sense locality, providing a comprehensive assessment of LMs’ adaptability and robustness under dynamic knowledge changes. The differences between CRAFT and previous datasets are summarized in Table 1.
We perform exposure analysis on CRAFT and other evaluation datasets with five different LMs. The results indicate that a large portion of knowledge in existing datasets has been known to LMs while the exposure rate of CRAFT is nearly zero. This validates the real-time property of CRAFT, which makes it challenging and fair.
We evaluate a suite of representative knowledge editing methods on CRAFT and demonstrate current approaches exhibit inherent limitations. Parameter-based approaches like ROME Meng et al. (2022) and WISE Wang et al. (2024a) suffer from gradual model degradation and scarcely edit successfully when new knowledge comes sequentially and continually. Simple retrieval-based approaches like IKE Zheng et al. (2023) and EREN Chen et al. (2024b) fail to achieve consistent balanced performance due to lack of alignment. LTE Jiang et al. (2024), which aligns LMs with knowledge editing via finetuning, shows weak locality due to overfitting on irrelevant queries.
@@
  • We reveal the leakage issue of existing datasets and introduce CRAFT, an open-source, contamination-free and easily-updated dataset for real-time knowledge editing.
  • We evaluate existing knowledge editing methods on CRAFT and reveal the limitations of these methods in real-time knowledge editing, especially the tradeoff among edit success, locality and portability.
  • We propose KEDAS, a novel paradigm of aligning LMs with real-time knowledge editing which significantly outperforms existing methods with balanced performance.
This paper is organized as follows. We introduce the preliminary of knowledge editing in § 2. Then, we present our CRAFT dataset in § 3 and propose KEDAS for real-time knowledge editing in § 4. After that, we conduct empirical experiments in § 5. Finally, we discuss related work in § 6 and conclude our study in § 7.
@@ 2.1 Task Formulation
An LM system $f$ can be regarded as a function $f:\mathcal{Q}\mapsto\mathcal{A}$ mapping an query $q\in\mathcal{Q}$ to its output answer $a=f(q)\in\mathcal{A}$.
Knowledge editing aims to change the behavior of the LM after modifying some knowledge. Edits of knowledge are usually in the form of query-answer (QA) pairs:
@@
To assess the efficacy of knowledge editing, the post-edit LM $f^{*}$ is evaluated via the following criteria Zhang et al. (2024):
@@ 2.2 Settings of Knowledge Editing
This traditional setting is widely employed by earlier work. It is far from real-world applications, since only one edit can be applied.
@@
In sequential editing, edits are applied step by step in a continual manner ($f^{0}$ denotes $f$ for consistency):
@@
2.3 Knowledge Editing Alignment
LTE Jiang et al. (2024) relates knowledge editing to LM alignment, aligning LMs with ever-changing real-time knowledge edits. During alignment, LMs’ capabilities of knowledge updating are enlightened by an editing prompt:
@@ 2.3 Knowledge Editing Alignment
In the inference phase, LMs are required to conduct on-the-fly and streaming knowledge editing by retrieving relevant edits to the query from the stored memory.
@@
MQuAKE-T Zhong et al. (2023) 35.46% 76.21% 46.11% 42.58% 76.09% 5,604
[…]
@@ 3 The CRAFT Dataset
We introduce CRAFT, a real-time and real-world dataset designed to evaluate knowledge editing under dynamically evolving factual contexts. Unlike static datasets that rely on outdated or widely exposed Wikipedia facts, CRAFT continuously updates with publicly available official data sources, ensuring fairness and real-world relevance.
3.1 Data Sources
CRAFT consists of two complementary subsets, namely CRAFT-Statistics and CRAFT-Finance, representing two highly correlated real-world domains that evolve periodically. Example instances from the two subsets are shown in Appendix G and H respectively.
@@ 3.1 Data Sources
We collect monthly statistical reports from the National Bureau of Statistics of China [^3] https://data.stats.gov.cn/, covering July 2024 to June 2025 in the current version used in this work. The dataset spans four major domains – finance, fiscal affairs, telecommunications, and transportation – containing 221 indicators and 4,468 data points. Our data collection pipeline, leveraging the cn-stats API [^4] https://github.com/songjian/cnstats, supports automatic temporal updates, enabling effortless collection of the most recent monthly data.
@@
CRAFT-Finance includes annual financial statements of 390 publicly listed Chinese companies (2023–2024), collected via the AKShare API [^5] https://github.com/akfamily/akshare. It features financial indicators such as debt ratio, shareholder equity, and operating cash flow, supporting yearly updates to reflect the latest financial disclosures.
@@ 3.2 Evaluation Metrics
To comprehensively assess model adaptability, CRAFT provides four evaluation dimensions besides the conventional edit success (ES) as described in § 2.1:
@@
$q_{p}$ $:\text{2024 Company A's debt ratio}$
$a_{p}$ $:0.5$
CRAFT-Finance includes 10 types of such combinations, demonstrating the generality and practical relevance of composite reasoning portability. For CRAFT-Statistics, composite portability is constructed by pairing two edits of the same indicator from adjacent years.
@@
3.3 Exposure Analysis
A perfect evaluation dataset should be unleaked, with a low rate of knowledge exposed to LMs (i.e., exposure rate). If a large portion of knowledge in a knowledge editing dataset has already been known by LMs, the evaluation results on it can be inaccurate, unfair, or even misleading.
We evaluate the exposure rates of CRAFT and existing benchmarks across five representative LMs by directly prompting with the questions in the datasets and counting how many questions can be correctly answered. As shown in Table 2, Due to its temporal freshness and domain specificity, CRAFT exhibits extremely low exposure rates (nearly 0%) across all the five LMs, mitigating data contamination and ensuring fair evaluation. Meanwhile, for other datasets, a great portion of the knowledge has been known to the LMs, making them less challenging and less fair in practice. This reveals the advantage of CRAFT as a truly real-time dataset.
@@ 3.4 Dataset Statistics
Subset Edit $\mathbf{P}_{\text{reasoning}}$ $\mathbf{P}_{\text{alias}}$ $\mathbf{L}_{\text{temporal}}$ $\mathbf{L}_{\text{common}}$
@@
Table 3: The numbers of instances in different evaluation types of CRAFT. Composite reasoning uses edit pairs; other metrics are per edit.
Table 3 shows the numbers of instances in CRAFT’s subsets. Composite reasoning portability is evaluated per edit pair, and other metrics per edit. For experiments, we split each subset of CRAFT into training and test sets. The test set contains 500 instances from the two subsets each (1,000 instances in total) and the training set contains the remaining instances.
3.5 Discussion
CRAFT is a scalable and continuously evolving dataset reflecting real-world knowledge updates. Its dynamic nature, factual authenticity, and comprehensive reasoning evaluation make it a robust foundation for future research on reliable knowledge editing. Anyone who wants to construct CRAFT within an arbitrary time period can easily obtain a customized dataset with our provided open-source scripts. Through CRAFT, we also demonstrate that researchers can make full use of public available data sources (like public APIs) besides the widely used Wikipedia data to create datasets with real-world and real-time properties and scalability that can help evaluate, train or develop more scalable intelligent systems without much human labor.
@@ 4 KEDAS: A Proposal for Real-time Knowledge Editing Alignment
As will be shown in the experiments (§ 5), previous methods show certain limitations either in recall (i.e., edit success and portability) or precision (i.e., locality) on CRAFT. To improve edit success and generalization while keeping a good locality in real-time knowledge editing, we propose KEDAS, an all-round approach consisting of several complementary designs.
The overall framework of KEDAS is illustrated in Figure 1, including alignment, editing and inference. In alignment, the LM is aligned with knowledge editing using LoRA. In editing, each edit is converted into diverse forms and stored to the memory module. In the inference phase, edits are retrieved from the memory based on the user query and then filtered by a trained classifier. If there exist relevant edits, the model with trained LoRA adapters will be called for inference. Otherwise, the query will go through the frozen base model. The order of editing and inference is flexible based on practical needs (“edit as you go”), as shown by the bidirectional arrow. For example, we may apply several edits and employ the LM for inference immediately. If there come further new edits, we can return to the editing phase to perform these edits and then resume inference.
@@ 4.2 Diverse Edit Augmentation
In previous work, each piece of edited knowledge is directly applied to the system as it is, limiting generalization. Since each knowledge can be expressed in various ways, leveraging just one fixed form is inflexible, especially for challenging portability evaluations like that in CRAFT.
@@
Concretely, for an edit $e^{t}=(q^{t}_{e},a^{t}_{e})$ at step tt [^6] For simplicity, we deem the two edits in each CRAFT instance as two separate time steps., its conventional QA form is:
@@
Specifically, declarative is to convert the QA pair into a declaration while aliased is to replace specialized terms in the declaration with their aliases for more flexible retrieval. In experiments, we conduct these augmentations by prompting an API-based off-the-shelf LM.
@@ 4.3 Filter-based Smart Retriever
During inference, the common practice regarding retrieval in previous knowledge editing work is to directly employ an embedder and perform single-stage retrieval to obtain the top-$k$ hits. This suffers from a poor locality due to the lack of a filter measuring the relevance of queries. In this work, we propose a filter-based smart retriever that pairs the high-recall retrieval with a high-precision filter by precisely identifying relevant queries. In this way, on challenging datasets like CRAFT, the post-edit system can keep consistent behaviors for neighboring locality queries and preserve its original answers to general queries.
First, we retrieve top-$n$ candidates from the memory using a normal embedding-based retriever. Thanks to our diverse edit augmentation, the memory contains various forms of edits, improving the recall rate. At the time step $t$, for an input query $q$, a set of top-$n$ candidates $\mathcal{C}^{\text{top-}n}$ is retrieved as:
@@
$\mathcal{E}^{*}=\operatorname{arg\,top}_{k}(\mathbf{Sim}(e,q))\ \text{for }e\in\mathcal{C}^{\text{filtered}}.$ (13)